-> 6A
+CB
---------
1B7
In the correctly worked addition problem above, each of the three letters A,B,C represents a different digit and B+C = 14. What is the value of A then?

Answer by CubeyThePenguin(3113)

....6A
+...CB
-------
..1B7

(60 + A) + (10C + B) = 100 + 10B + 7

A - 9B + 10C = 47
9B + 9C = 126

A + 19C = 173

173/19 = 9 = C and A = 2